hi i would like your help to help me with this question. Thanks
quesion: 1/r=1/r1 + 1/r2
Express r1 in terms of r and r2

Answer by mananth(16949)

1/r=1/r1 + 1/r2
add -1/r2 to both sides
1/r - 1/r2 = 1/r1
r2-r / rr2= 1/r1
r1=rr2/r2-r